A total of 2. 00 mol of a compound is allowed to r
Answer #1

The  enthalpy of reaction is given by

∆H = m x s x ∆T

m = mass = 189 g

s = specific heat = 4.179 J/ g oC (water)

∆T = Change in temperature = 24.70-21.00= 3.7 oC

∆H =  189 * 4.179 * 3.7 = 2922.375 J

This value is for two moles of compound.

The enthalpy of reaction = 2922.375 /2 = 1461.187 J/mol